Previous

Calendar

July

2009

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

28

JD179

29

JD180
Site Visit (SpB, CG, AJ, SO, L): Details , Photos (25)

Span cylinder at 1800 psi. N2 at 200 psi --> 2000 psi.

at ~10:30 MST, reconnected channel 200 (UC logger); added in-line surge protector at tower end.

from 13:00-14:10 MST, switched N2 cylinder at the tower (new cylinder=CC199722 (2000 psi); old cylinder=CC198519 (200 psi))

from 12:00-12:30 MST, made more adjustments to h2o.ec offset using dial (see details)

from 13:20-13:45 MST, replaced domes on REBS Rnet sensor

TRAM was running!

Hot and Dry. Cu clouds formed as the day went on. It rained briefly at around 13:30 MST.

30

JD181

1

JD182

2

JD183

3

JD184
CU Holiday?

4

JD185
Independence Day

5

JD186

6

JD187
Site Visit (SpB, LA, CG, AJ): Details , Photos (12)

Span cylinder at 1710 psi. N2 at 1760 psi.

mucked around with network to TRAM Adam (still not working)

Cool and Dry. High clouds, but clouded up in the late afternoon.

7

JD188
Site Visit (SpB, LT, SL, LA): Details , Photos (18)

at ~10:00 MST, added KZ PAR-LITE sensor as PAR2.up to CNR logger (note there is a 4 deg N-S tilt)

Did a "practice" re-calibration of the 2.5m LI-7500 (see notes)

West Robot removed (Steve R.)

Starting to warm up. Primarily clear blue skies all day (no rain).

(Full "Buck" Moon)

8

JD189

9

JD190

10

JD191
Site Visit (SpB, LA, CG, AJ, LL, R, +): Details , Photos (13)

swapped LI-6262 out for re-cal (replaced with LI-6262 loaner, sn IRG3-224)

changed range for h2o.ec back to 1000 mV on "fast" cr23x logger (fast_09.csi).

Warm.

11

JD192

12

JD193

13

JD194
Site Visit (SpB, CG, AJ, SL, NT, +): Details , Photos (11)

Span cylinder at 1700 psi. N2 at 1600 psi.

installed MATLAB on laptops in the computer classroom

A brief T-storm passed by tower at 13:00...cooled things off.

Start of the NSF "Flux" class at MRS.

14

JD195

15

JD196
Site Visit (SpB, +): Details

Flux class viewed campbell "demonstration" setup at C1.

16

JD197
Site Visit (SpB, CG, AJ, SL, SO, +): Details , Photos (7)

Flux class visited the AmeriFlux tower; S. Oncley also giving a site-tour.

Clear blue, sunny, and HOT.

17

JD198
Site Visit (SpB, +): Details

Morning Lectures by D.Bowling and D.Baldocchi

18

JD199

19

JD200

20

JD201

21

JD202

22

JD203
Site Visit (SpB, LA, CG, AJ, +): Details , Photos (25)

Span cylinder at 1600 psi. N2 at 1350 psi.

at 12:00 MST, leveled the PAR-LITE sensor

from 12:06-12:30 MST, Changed profiler (li-6251) dessicant

added a 5-port switch at the 10-m power box to prepare for the new webcam

unplugged the heater for the rain-gauge

Big puffy clouds.

23

JD204
Site Visit (SpB, LA, AR, MC, CG, AJ, +): Details , Photos (13)

added a new "StarDot" netcam XL webcam, e.g.: Image

Some clouds around. A brief shower and dark clouds in the afternoon. Followed by clear blue skies.

24

JD205
End of the NSF "Flux" class at MRS.

25

JD206

26

JD207

27

JD208

28

JD209
Report Jeep Mileage to Lease Operations

Heavy Fog/Overcast.

29

JD210
Site Visit (SpB, AJ): Details , Photos (29)

Span cylinder at 1500 psi. N2 at 1120 psi.

Added a T/RH sensor from EOL (currently setup at 188 cm) to prof logger (start using prf_02.csi)

Added SDM output from the LI-7500 at 2.5m, uc23x logger starts using uc_02.csi (note: this change causes data output rate to be at 5 Hz, not 10 Hz)

Foggy/overcast with some rain.

30

JD211
(note: csat at 21.5m was not working most of the day due to fog/rain)

at 9:10 MST, started ops50; new LI-7500 parameters from uc23x + T ("o3.slow") and RH (RH.eol) from eol T/RH sensor

Foggy.

31

JD212
Jeep gets a flat tire (LL)

1

JD213